Discrete math Show that for k=2 n(n-1) (n chooses k) 5^k = 25n(n-1)6^(n-2) combinatorically.

Discrete math

Show that for k=2 n(n-1) (n chooses k) 5^k = 25n(n-1)6^(n-2) combinatorically.
Answer #1

men ner 2 -5 5k=n(n-1) Taking 5= n(n-1) EGM 7 For K=2, n(n-1) sk= 25n(n) (0-2) n! n! -H.S=> niny) 6) K! M-2) 12 n(n-1) (2) an2 Putting n=2 na2 in both sides Sa So, L.H.Say 25m² (n.12 25(3321213214,3 = 25x2 =50 25 nin- = 251296272 62 25x2 = 50 62 non)